# Energy-Efficient Transformer Inference: Optimization Strategies for Time Series Classification

### v2

@@ Abstract
The increasing computational demands of transformer models in time series classification necessitate effective optimization strategies for energy-efficient deployment. This paper presents a systematic investigation of optimization techniques, focusing on structured pruning and quantization methods for transformer architectures. Through extensive experimentation on three distinct datasets (RefrigerationDevices, ElectricDevices, and PLAID), we quantitatively evaluate model performance and energy efficiency across different transformer configurations. Our experimental results demonstrate that static quantization reduces energy consumption by 29.14% while maintaining classification performance, and L1 pruning achieves a 63% improvement in inference speed with minimal accuracy degradation. These findings provide valuable insights into the effectiveness of optimization strategies for transformer-based time series classification, establishing a foundation for efficient model deployment in resource-constrained environments.
@@ Introduction
Recent advancements in transformer architectures have significantly advanced time series analysis across various domains, including healthcare, finance, and predictive maintenance, enabling more accurate forecasting, anomaly detection, and decision-making processes [8]. The ability of AI models to process sequential data efficiently has enabled substantial improvements in these fields, allowing for more accurate forecasting, anomaly detection, and decision-making processes [23]. At the core of these developments are transformer-based architectures, which have demonstrated superior performance over traditional models, such as recurrent neural networks (RNNs) and long short-term memory (LSTM) networks, by leveraging self-attention mechanisms to capture long-range dependencies in temporal data [29]. Despite their effectiveness, transformers are computationally expensive, making them less viable for real-time and edge-based applications due to their high energy consumption and memory footprint.
The growing carbon footprint associated with transformer training and inference has raised significant concerns regarding sustainability and deployment feasibility in resource-constrained environments [20]. As AI adoption scales across industries, the carbon footprint of large-scale models has become a pressing issue, necessitating the development of optimization strategies that reduce computational overhead while maintaining predictive accuracy.
@@
Transformers utilize self-attention mechanisms to capture long-range dependencies in sequential data, making them particularly effective for time series forecasting and classification [25]. Traditional models, such as recurrent neural networks (RNNs) and long short-term memory (LSTM) networks, struggle with long-range dependencies due to their sequential nature, often suffering from vanishing gradient issues. In contrast, transformers process all time steps simultaneously, enhancing their ability to learn complex temporal patterns [31].
@@
Domain-specific preprocessing has further enhanced transformer performance. Wavelet transforms have been employed to enable multi-resolution representations of time series data, improving their ability to capture both local and global patterns [21]. Additionally, ensemble learning techniques have been explored to enhance forecasting accuracy [16].
Empirical evaluations have reinforced transformers’ effectiveness in time series forecasting and classification. A study by Lara-Benítez et al. [15] analyzed transformer performance across 12 datasets with over 50,000 time series, providing crucial insights into their advantages and limitations across various domains.
Model Compression Strategies: Pruning and Quantization
To mitigate computational overhead, several model compression strategies have been proposed, including pruning, quantization, and knowledge distillation.
Pruning is a well-established method for reducing model size by removing parameters deemed less significant. This technique has been successfully applied to transformers to improve efficiency while preserving accuracy. Pruning-guided feature distillation has been introduced to create lightweight transformer architectures that maintain predictive performance while reducing computational costs [12]. Additionally, global structural pruning has demonstrated significant reductions in latency and computational requirements [18]. Cheong [3] further highlights the role of pruning in compressing transformers to enhance inference speed and energy efficiency.
Quantization, which reduces model weight and activation precision, is another widely used approach for decreasing memory usage and improving inference speed. Quantization-aware training has shown promise in minimizing memory footprints while preserving accuracy [32]. Techniques such as mixed-precision quantization [27], post-training quantization [18], and quantized feature distillation have been effective in reducing resource consumption.
Knowledge distillation, which trains a smaller model to replicate the performance of a larger one, further contributes to reducing model complexity while maintaining accuracy. This approach has proven particularly useful for deploying transformers in resource-constrained environments [28]
@@
The NIOT framework, specifically designed for modern CPUs, integrates architecture-aware optimizations such as memory tiling, thread allocation, and cache-friendly fusion strategies. These improvements have led to latency reductions of up to 29% for BERT and 43% for vision transformers, significantly surpassing traditional inference techniques [30].
@@
The sustainability of transformer models has been analyzed within the broader framework of green computing.  [14] introduce the concept of green algorithms, providing a quantitative framework for assessing the carbon footprint of computational tasks. This metric is instrumental in evaluating the environmental impact of transformer-based architectures in time series classification, reinforcing the importance of computational efficiency in sustainable AI practices.
@@
Several studies have examined the trade-offs between performance and energy efficiency in transformer inference. [2, 7] present empirical evaluations illustrating how software-level optimizations can significantly lower energy consumption without sacrificing predictive accuracy. These findings underscore the necessity for targeted optimization strategies, particularly for CPU-based inference, where resource constraints are a fundamental challenge.
@@
Complementary to these efforts,  [1] introduce the Greenup, Powerup, and Speedup (GPS-UP) metrics to evaluate energy efficiency in software optimizations. By categorizing computational trade-offs into multiple distinct scenarios, their study provides a structured framework for analyzing the relationship between software modifications and energy consumption. Unlike conventional energy-delay metrics, GPS-UP facilitates a more nuanced evaluation of how performance improvements interact with power efficiency, contributing to the development of sustainable yet high-performance transformer models.
Background
Time Series Signal Processing
Time series signals in real-world applications can be represented as a sequence $\mathbf{X}=\{x_{t}\}_{t=1}^{T}$, where each $x_{t}\in\mathbb{R}^{d}$ represents a $d$-dimensional measurement at time step $t$. Traditional approaches to processing such signals have relied on various architectures, each with distinct characteristics in capturing temporal dependencies. Consider a time series signal $x(t)$; its discrete-time representation can be expressed as:
$x[n]=x(nT_{s}),\quad n\in\mathbb{Z}$ (1)
where $T_{s}$ is the sampling period. The corresponding frequency domain representation through the Discrete Fourier Transform (DFT) is:
$X[k]=\sum_{n=0}^{N-1}x[n]e^{-j2\pi kn/N}$ (2)
Deep Learning for Sequential Data
Deep learning approaches for sequential data processing have evolved from recurrent architectures to attention-based mechanisms. The traditional recurrent neural network (RNN) processes sequential data through:
$h_{t}=\sigma(W_{xh}x_{t}+W_{hh}h_{t-1}+b_{h})$ (3)
where $h_{t}$ represents the hidden state at time $t$, $W_{xh}$ and $W_{hh}$ are weight matrices, and $\sigma(\cdot)$ is a nonlinear activation function. While RNNs are effective at processing sequences, they face challenges when trying to capture relationships between elements that are far apart in the sequence. This difficulty is largely due to the vanishing gradient problem, a phenomenon where the influence of earlier inputs diminishes as the sequence progresses. As a result, RNNs can struggle to learn long-range dependencies, meaning they might not effectively remember or use information from many time steps earlier in the sequence.
To overcome these limitations, researchers have developed more advanced architectures. One of the most significant advancements has been the introduction of attention-based mechanisms, allowing models to focus on specific parts of the input sequence, regardless of their position. This led to more powerful models that can handle long-range dependencies more effectively than traditional RNNs.
@@
Optimizations for Energy Efficiency
Enhancing the efficiency of Transformer-based models necessitates the use of optimization strategies, such as quantization, pruning, and specialized hardware acceleration. These techniques aim to reduce computational overhead, lower energy consumption, and improve inference speed without significantly compromising model accuracy.
Quantization
Quantization is a widely adopted method for reducing computational complexity and memory requirements by representing model parameters and activations with lower precision data types. Typically, deep learning models are trained using high-precision 32-bit floating-point (FP32) representations; however, during inference, these representations can be converted to lower bit-width formats, such as 8-bit integers (INT8), significantly reducing memory bandwidth and energy consumption.
$E_{quantized}=\frac{E_{float32}}{Q}$ (6)
where $Q$ denotes the quantization factor. Reducing numerical precision can lead to substantial power savings while maintaining a minimal impact on model performance.
Pruning
Pruning is an effective strategy for reducing computational complexity by eliminating redundant parameters that contribute minimally to the model’s predictive performance. Transformer architectures often contain a significant number of parameters, many of which are unnecessary for effective inference.
A variety of both structured and unstructured pruning techniques have been developed to optimize Transformers. These methods include weight pruning, which involves removing individual weights that have minimal impact on the model’s accuracy; neuron pruning, where entire neurons in dense layers are discarded to reduce computational cost; and head pruning, which targets and eliminates redundant attention heads within the multi-head self-attention mechanisms. Additionally, structured pruning can remove entire layers or blocks from the network, thereby significantly enhancing overall efficiency. The impact of pruning on the model’s complexity can be quantified by
$E_{pruned}=E\times(1-p)$ (7)
where $p$ denotes the proportion of parameters removed.
When pruning is combined with quantization, the benefits become even more pronounced, as pruned models require fewer computations, while quantized models utilize lower-bit representations, leading to significant reductions in power consumption and memory bandwidth usage.
Methodology
This study aims to optimize transformer-based models for energy-efficient time series classification. The methodology consists of structured data preprocessing, an optimized transformer architecture, and energy-efficient optimization strategies such as pruning and quantization. Three datasets—RefrigerationDevices, ElectricDevices, and PLAID—are employed for evaluation. The experimental setup involves training transformer models on these datasets and measuring their computational efficiency, classification performance, and power consumption.
Dataset Preprocessing
The datasets used in this study contain multivariate time series data from various electrical devices. To ensure consistency across datasets and improve the performance of the transformer models, a structured preprocessing pipeline is applied:
Segmentation: Fixed-length overlapping windows are extracted from each time series to ensure that sufficient contextual information is available for classification.
Train-Validation-Test Splitting: A subject-wise splitting strategy is used, ensuring that data from the same household does not appear in both training and test sets, preventing data leakage.
Transformer Model Implementation
This study incorporates the Vision Transformer (ViT) model [5], which has been increasingly adapted for time series tasks due to its ability to capture long-range dependencies efficiently [11]. The model architecture comprises several key components. First, the Time Series Patch Embedding Layer transforms raw time series data into fixed-size patches using a one-dimensional convolutional layer, with positional encoding added to maintain the sequential order. Next, the multi-head self-attention mechanism allows the model to focus on different segments of the time series simultaneously, enhancing its ability to learn complex relationships. Each encoder block contains position-wise feed-forward layers with ReLU activations and dropout layers to reduce overfitting. To ensure stable training and mitigate the risk of vanishing gradients, the architecture incorporates layer normalization and residual connections. Finally, a fully connected softmax output layer classifies each sequence into its corresponding device category.
Optimization Strategies
To enhance the computational efficiency and energy sustainability of transformer models for time series classification, two primary optimization techniques—pruning and quantization—are implemented.
5.3.1 Pruning
In this research, two main pruning approaches are explored. The first approach is Magnitude-Based Pruning, which eliminates parameters with the smallest absolute values under the assumption that lower-magnitude weights contribute less to overall model performance. In this context, L1-Norm Pruning removes weights with the smallest L1 norm, thereby reducing network sparsity while largely preserving accuracy. Alternatively, L2-Norm Pruning removes entire neurons or filters based on their L2 norm (Euclidean distance), effectively eliminating redundant units while maintaining structural integrity. Additionally, Structured Pruning is applied to remove entire layers, filters, or attention heads, thereby significantly reducing overall model complexity.
5.3.2 Quantization
In our study, we employ two distinct quantization strategies to decrease both memory footprint and computational overhead while maintaining high model accuracy.
First, we apply Post-Training Quantization (PTQ). It reduces the precision of model weights and activations after training by converting 32-bit floating-point numbers into lower-bit representations (e.g., 8-bit or 16-bit). This process significantly decreases the model’s memory requirements and computational load. Within PTQ, we implement two approaches. In static quantization, model weights and activations are converted to lower precision prior to inference, whereas in dynamic quantization, only the weights are quantized while activations remain in floating-point representation.
Our implementation of quantization is carried out using the torch.quantization module in PyTorch. The quantization workflow involves three main stages: model preparation, quantization configuration, and calibration with evaluation. During model preparation, we select specific layers—such as linear projections and self-attention heads—that are critical for performance and amenable to quantization. In the quantization configuration step, high-precision floating-point values are replaced with integer-based representations (e.g., int8). Finally, we perform post-quantization calibration using validation data to ensure that any loss in accuracy is kept to a minimum.
[…]
@@ 5.3.2 Quantization
Experiments are conducted in a GPU-accelerated computing environment using the PyTorch framework along with the TorchQuantization library. To ensure statistical robustness, each model configuration is trained over multiple runs, and the resulting performance metrics are averaged. This approach guarantees that the reported results are reliable and reflective of the models’ true performance across diverse conditions.
Training Strategy and Hyperparameters
In our experiments, we adopted a carefully designed training strategy to ensure robust model convergence while optimizing computational resources. Specifically, we set the batch size to 64 samples per batch, a choice made to optimize memory usage and facilitate efficient gradient updates. The model parameters were optimized using the Adam optimizer, enhanced with weight decay regularization to mitigate overfitting.
A cosine annealing scheduler was employed to adaptively adjust the learning rate during training, thereby promoting a smooth convergence process. For multi-class classification tasks, the categorical cross-entropy loss function was utilized to quantify prediction errors. The training was conducted over 50 epochs, with early stopping based on validation accuracy to prevent overfitting and to conserve computational resources.
Furthermore, model checkpoints were systematically saved throughout the training process, and the best-performing model was ultimately selected based on the minimum validation loss observed.
@@
In our study, we employ three distinct datasets to analyze and classify electrical device usage patterns: RefrigerationDevices, ElectricDevices, and PLAID.
RefrigerationDevices Dataset The RefrigerationDevices dataset is part of the UCR Time Series Classification Archive [6]. It comprises time series data specifically focusing on household refrigeration appliances. Each time series consists of 720 data points, representing 24 hours of readings taken at two-minute intervals. This granularity facilitates an in-depth analysis of daily usage patterns of refrigeration devices. The dataset’s focus on refrigeration appliances makes it particularly relevant for studies aiming to understand and optimize energy consumption in household refrigeration, a significant component of residential energy use.
ElectricDevices Dataset Also sourced from the UCR Time Series Classification Archive [6], the ElectricDevices dataset encompasses a broader range of household appliances beyond refrigeration. It includes data from various electrical devices used in homes, divided into a training set with 8,926 instances and a test set with 7,711 instances. Each time series comprises 96 data points, capturing the operational characteristics of different devices across seven distinct categories. This comprehensive dataset allows for the analysis of diverse appliance usage patterns, providing insights into the classification and energy consumption of various household devices.
PLAID Dataset The Plug Load Appliance Identification Dataset (PLAID) is designed for load identification research. It includes short voltage and current measurements sampled at 30 kHz from 11 different appliance types present in more than 60 households in Pittsburgh, Pennsylvania, USA. Data collection occurred during the summer of 2013 and winter of 2014. Each appliance type is represented by multiple instances of varying make and models, with three to six measurements collected for each state transition. These measurements were post-processed to extract windows containing both steady-state operation and startup transients. After removing measurements with significant noise due to errors, the dataset comprises 1,074 instances. The high sampling rate and detailed measurements make PLAID particularly valuable for developing and testing algorithms aimed at non-intrusive load monitoring and appliance identification.
@@
Experimental Setup
In order to comprehensively evaluate the effectiveness of our proposed optimization framework, we conducted extensive experiments utilizing three distinct datasets that represent varying aspects of electrical device usage patterns and characteristics. The RefrigerationDevices dataset, comprising 2,500 sequences with 720 timesteps each, provides a focused examination of temporal patterns in refrigeration systems, while the ElectricDevices dataset, with its substantial training set of 8,926 instances and testing set of 7,711 instances, offers a broader perspective on general electrical device behavior. Additionally, the PLAID dataset, containing 1,074 instances distributed across 11 distinct device categories, enables the evaluation of our framework’s generalization capabilities across diverse appliance types and usage patterns.
Implementation Details and Model Configurations
Performance Analysis
7.3.1 Classification Accuracy Assessment
@@
7.3.2 Impact of Optimization Techniques

### v3

@@ Abstract
The increasing computational demands of transformer models in time series classification necessitate effective optimization strategies for energy-efficient deployment. Our study presents a systematic investigation of optimization techniques, focusing on structured pruning and quantization methods for transformer architectures. Through extensive experimentation on three distinct datasets (RefrigerationDevices, ElectricDevices, and PLAID), we quantitatively evaluate model performance and energy efficiency across different transformer configurations. Our experimental results demonstrate that static quantization reduces energy consumption by 29.14% while maintaining classification performance, and L1 pruning achieves a 63% improvement in inference speed with minimal accuracy degradation. Our findings provide valuable insights into the effectiveness of optimization strategies for transformer-based time series classification, establishing a foundation for efficient model deployment in resource-constrained environments.
@@ Introduction
Recent advancements in transformer architectures have significantly advanced time series analysis across various domains, including healthcare, finance, and predictive maintenance, enabling more accurate forecasting, anomaly detection, and decision-making processes [9]. The ability of AI models to process sequential data efficiently has enabled substantial improvements in these fields, allowing for more accurate forecasting, anomaly detection, and decision-making processes [25]. At the core of these developments are transformer-based architectures, which have demonstrated superior performance over traditional models, such as recurrent neural networks (RNNs) and long short-term memory (LSTM) networks, by leveraging self-attention mechanisms to capture long-range dependencies in temporal data [31]. Despite their effectiveness, transformers are computationally expensive, making them less viable for real-time and edge-based applications due to their high energy consumption and memory footprint. Furthermore, the growing carbon footprint associated with transformer training and inference has raised significant concerns regarding sustainability and deployment feasibility in resource-constrained environments [22]. With the exponentially increasing usage of AI, the carbon footprint of massive models has been a topic of increasing worry. At the same time, deep learning model’s runaway scaling, such as huge transformers and diffusion models, has induced unmatched computation costs, which are enormous and require a lot of power to operate and train, directly contributing to the increase in global carbon emissions. Consequently, feeding environmental concerns such as global warming. If not checked, the long-term growth of AI systems will hugely frustrate efforts at sustainability and drain energy resources.
focusing to encompass computation overheads, researchers are developing strategies not to compromise forecasting performance. Enhancing model efficiency, in addition to techniques like knowledge distillation, pruning, quantization, and even hardware-based architectures with increased energy efficiency, should be a pursuit by research professionals. Furthermore, green AI technologies, i.e., renewable energy for data centers and minimizing the algorithms to consume less energy, need to be transitioned in order to make AI a force for good and not an environmental issue.
@@
Transformers utilize self-attention mechanisms to capture long-range dependencies in sequential data, making them particularly effective for time series forecasting and classification [28, 26]. Traditional models, such as recurrent neural networks (RNNs) and long short-term memory (LSTM) networks, struggle with long-range dependencies due to their sequential nature, often suffering from vanishing gradient issues. In contrast, transformers process all time steps simultaneously, enhancing their ability to learn complex temporal patterns [34].
@@
Domain-specific preprocessing has further enhanced transformer performance. Wavelet transforms have been employed to enable multi-resolution representations of time series data, improving their ability to capture both local and global patterns [23]. In addition, ensemble learning techniques have been explored to enhance forecasting accuracy [18].
Empirical evaluations have reinforced transformers’ effectiveness in time series forecasting and classification. A study by Lara-Benítez et al. [17] analyzed transformer performance across 12 datasets, providing crucial insights into their advantages and limitations across various domains. This comprehensive evaluation not only highlighted the strong performance of transformers in various settings but also revealed specific scenarios where they perform exceptionally well, as well as cases where their effectiveness might be limited.
Model Compression Strategies
In recent years, model compression has gained attention in the domain of deep learning as it directly addresses challenges related to the large size and computational demands of neural networks. Two prominent strategies employed in model compression are pruning and quantization.
Pruning is a well-established method for reducing model size by removing parameters deemed less significant. This technique has been applied to transformers to improve efficiency while preserving accuracy. Pruning-guided feature distillation has been introduced to create lightweight transformer architectures that maintain predictive performance while reducing computational costs [14]. Additionally, global structural pruning has demonstrated significant reductions in latency and computational requirements [20]. Cheong [3] further highlights the role of pruning in compressing transformers to enhance inference speed and energy efficiency.
Quantization reduces the precision of model weights and activations, making it a popular method for decreasing memory usage and enhancing inference speed. Quantization-aware training has demonstrated effectiveness in reducing memory footprints while maintaining accuracy [35]. Techniques such as mixed-precision quantization [30], post-training quantization [20], and quantized feature distillation have been effective in reducing resource consumption.
@@
A study utilized the NIOT framework, specifically designed for modern CPUs, which integrates architecture-aware optimizations such as memory tiling, thread allocation, and cache-friendly fusion strategies. These improvements have led to latency reductions of up to 29% for BERT [7] and 43% for vision transformers, significantly surpassing traditional inference techniques [33].
@@
The sustainability of transformer models has been analyzed within the broader framework of green computing. A study introduced the concept of green algorithms, providing a quantitative framework for assessing the carbon footprint of computational tasks [16]. This metric is instrumental in evaluating the environmental impact of transformer-based architectures in time series classification, reinforcing the importance of computational efficiency in sustainable AI practices.
@@
Several studies have examined the trade-offs between performance and energy efficiency in transformer inference. Some studies present empirical evaluations illustrating how software-level optimizations can significantly lower energy consumption without sacrificing predictive accuracy [2, 8]. These findings underscore the necessity for targeted optimization strategies, particularly for CPU-based inference, where resource constraints are a fundamental challenge.
@@
In addition to these studies, a research introduced the Greenup, Powerup, and Speedup (GPS-UP) metrics to evaluate energy efficiency in software optimizations  [1]. By categorizing computational trade-offs into multiple distinct scenarios, their study provides a structured framework for analyzing the relationship between software modifications and energy consumption. Unlike conventional energy-delay metrics, GPS-UP facilitates a more nuanced evaluation of how performance improvements interact with power efficiency, contributing to the development of sustainable yet high-performance transformer models.
Methodology
Our methodology consists of structured data preprocessing, an optimized transformer architecture, and energy-efficient optimization strategies, such as pruning and quantization. The methodological framework of this study comprises the following key steps:
Overview
In our approach, a continuous time series signal is modeled as
$\mathbf{X}=(\mathbf{x}_{1},\mathbf{x}_{2},\ldots,\mathbf{x}_{T}),$
where each observation $\mathbf{x}_{t}\in\mathbb{R}^{d}$ is a $d$-dimensional vector representing a measurement at time $t$. The objective is to generate a corresponding sequence of classification outputs
$\mathbf{Y}=(y_{1},y_{2},\ldots,y_{T}),$
with each label $y_{t}\in\{1,\ldots,K\}$ indicating the class assignment at the respective time step.
To effectively process the input signals, our methodology incorporates several key stages:
Data Preprocessing
The data utilized in our study contain multivariate time series data from various electrical devices. To ensure consistency across datasets and improve the performance of the transformer models, a structured preprocessing pipeline is applied:
Segmentation: The input time series is segmented into overlapping windows of fixed length $w$ with stride $s$. Let
$\mathbf{W}_{t}=(\mathbf{x}_{t-w+1},\mathbf{x}_{t-w+2},\ldots,\mathbf{x}_{t})$
denote the window ending at time $t$. Each window $\mathbf{W}_{t}$ is associated with a target label $y_{t}$. The complete dataset $\mathcal{D}$ consists of $N$ sequences, where each sequence contains multiple overlapping windows, ensuring that sufficient contextual information is available for classification.
Train-Validation-Test Splitting: A subject-wise splitting strategy is used, ensuring that data from the same household does not appear in both training and test sets, thereby preventing data leakage.
Transformer Model Implementation
Our study incorporates the Vision Transformer (ViT) model [5], which has been increasingly adapted for time series tasks due to its ability to capture long-range dependencies efficiently [13]. Our model architecture comprises several key components. First, the Time Series Patch Embedding Layer transforms raw time series data into fixed-size patches using a one-dimensional convolutional layer, with positional encoding added to maintain the sequential order. Next, the multi-head self-attention mechanism allows the model to focus on different segments of the time series simultaneously, enhancing its ability to learn complex relationships. Each encoder block contains position-wise feed-forward layers with ReLU activations and dropout layers to reduce overfitting. To ensure stable training and mitigate the risk of vanishing gradients, the architecture incorporates layer normalization and residual connections. Finally, a fully connected softmax output layer classifies each sequence into its corresponding device category.
Optimization Strategies
To enhance the computational efficiency and energy sustainability of transformer models for time series classification, pruning, and quantization are implemented.
3.4.1 Pruning
Transformer architectures often contain a significant number of parameters, many of which are unnecessary for effective inference. Pruning is an effective approach to reduce computational complexity by eliminating these redundant parameters that minimally contribute to the predictive performance of the model.
A variety of structured and unstructured pruning techniques have been developed to optimize Transformers. These methods include weight pruning [32], which involves removing individual weights that have minimal impact on the model’s accuracy; neuron pruning [11], where entire neurons in dense layers are discarded to reduce computational cost; and head pruning, which targets and eliminates redundant attention heads within the multi-head self-attention mechanisms. The impact of pruning on the model’s complexity can be quantified by
$E_{pruned}=E\times(1-p)$ (1)
where $p$ denotes the proportion of parameters removed.
In our research, two main pruning approaches are explored. The first approach is Magnitude-Based Pruning, which eliminates parameters with the smallest absolute values under the assumption that lower-magnitude weights contribute less to overall model performance. In this context, L1-Norm Pruning removes weights with the smallest L1 norm, thereby reducing network sparsity while largely preserving accuracy. Alternatively, L2-Norm Pruning removes entire neurons or filters based on their L2 norm (Euclidean distance), eliminating redundant units while maintaining structural integrity. Additionally, Structured Pruning is applied to remove entire layers, filters, or attention heads.
3.4.2 Quantization
Quantization is a widely adopted method to reduce computational complexity and memory requirement. Typically, deep learning models are trained using high-precision 32-bit floating-point (FP32) representations. However, during inference, these representations can be converted to lower bit-width formats, such as 8-bit integers (INT8).
$E_{quantized}=\frac{E_{float32}}{Q}$ (2)
where $Q$ denotes the quantization factor. Reducing numerical precision can lead to substantial power savings while maintaining a minimal impact on model performance. In our study, we employ two distinct quantization strategies to decrease memory footprint and computational overhead while maintaining high model accuracy.
First, we apply Post-Training Quantization (PTQ). It reduces the precision of model weights and activations after training by converting 32-bit floating-point numbers into lower-bit representations. This process significantly decreases the model’s memory requirements and computational load. Within PTQ, we implement two approaches. In static quantization, model weights and activations are converted to lower precision prior to inference, whereas in dynamic quantization, only the weights are quantized while activations remain in floating-point representation.
Our implementation of quantization is carried out using the torch.quantization module in PyTorch. The quantization workflow involves three main stage. We utilize model preparation, quantization configuration, and calibration with evaluation. During model preparation, we select specific layers—such as linear projections and self-attention heads—that are critical for performance and amenable to quantization. In the quantization configuration step, high-precision floating-point values are replaced with integer-based representations (e.g., int8). Finally, we perform post-quantization calibration using validation data to ensure that any loss in accuracy is kept to a minimum.
[…]
@@ 3.4.2 Quantization
Figure 3: Overall architecture
@@
Our experiments are conducted in a GPU-accelerated computing environment utilizing the PyTorch framework along with the TorchQuantization library. To ensure statistical robustness, each model configuration is trained over multiple runs, and the resulting performance metrics are averaged. Our approach guarantees that the results are reliable and reflective of the models’ performance across various conditions.
Implementation Details and Model Configurations
@@
Table 1: Statistics of the Time Series Datasets
Dataset Train Size Test Size Length  Classes
PLAID 537 537 30 11
ElectricDevices 8926 7711 96 7
RefrigerationDevices 375 375 720 3
In our study, we utilize three datasets to analyze and classify electrical device usage patterns: RefrigerationDevices, ElectricDevices, and PLAID.
RefrigerationDevices: From the UCR Archive [6], this dataset focuses on household refrigeration appliances, with each time series containing 720 data points (24 hours at two-minute intervals) for detailed daily analysis.
ElectricDevices: This dataset features 8,926 training and 7,711 test instances, each with 96 data points capturing appliance operations across seven categories. The goal was to gather data on household electricity usage behavior to aid in reducing the UK’s carbon emissions.
PLAID: Designed for load identification, PLAID includes high-frequency (30 kHz) voltage and current measurements from 11 appliance types across over 60 households. Collected in summer 2013 and winter 2014 and processed to extract relevant windows, it comprises 1,074 instances—ideal for non-intrusive load monitoring.
Table 1 summarizes their key properties, including training/test sizes, sequence lengths, and number of classes.
@@
Our experimental results reveal significant variations in model performance, with the PLAID dataset achieving the highest baseline accuracy (84.25% for T2). After optimization, models maintained high classification performance, with the lowest accuracy observed at 80.92% (after L2 pruning). In contrast, the RefrigerationDevices dataset, which had the lowest baseline accuracy (65.95% for T2), was more sensitive to optimization techniques, with accuracy dropping to 56.45% under L2 pruning.
The proposed framework demonstrates substantial improvements in energy efficiency, with static quantization achieving 29.14% energy savings while maintaining reasonable accuracy trade-offs. Notably, L1 pruning techniques achieved speed-ups of 1.63× compared to baseline models, while reducing energy consumption by 37.08%. The T2 architecture, despite higher computational complexity, provided superior accuracy-energy trade-offs across all datasets.
Classification Accuracy Assessment
@@
Impact of Optimization Techniques